Circular No. 6553

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

   URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1997 B1 (KOBAYASHI)

On Jan. 31 S. Nakano, Sumoto, Japan, reported observations on Jan. 30 and 31 by Takao Kobayashi of what appeared to be a minor planet. Further observations through Feb. 2 showed the orbit to be decidedly cometary. In response to our enquiry, W. Offutt verified that the object is indeed a comet. Available CCD precise positions follow, observers reporting the comet as asteroidal except as noted.

1997 UT α_{2000} δ_{2000} m_{2} Observer
Jan. 30.70931 10^{h}00^{m}20.^{s}77 +16^{o}29′55′′.6 18  Kobayashi
30.72829 10 00 20.04 +16 29 48.0  _{”}
31.66902 9 59 51.71 +16 22 47.8  _{”}
31.68528 9 59 51.20 +16 22 40.5  _{”}
Feb. 1.52586 9 59 25.23 +16 16 22.0  _{”}
1.77053 9 59 16.75 +16 14 31.5  _{”}
2.59105 9 58 50.59 +16 08 22.9 17.6 Abe
2.60067 9 58 50.28 +16 08 18.0  _{”}
2.83560 9 58 42.63 +16 06 28.4 17.7 Tichá
2.85667 9 58 41.92 +16 06 18.9  _{”}
3.20292 9 58 30.41 +16 03 43.9 17.8 Offutt
3.23778 9 58 29.17 +16 03 28.2 17.6  _{”}
3.56081 9 58 18.31 +16 01 00.5  Kojima
3.57109 9 58 17.97 +16 00 55.4 17.4  _{”}
3.90250 9 58 06.63 +15 58 23.2  Pravec
3.91046 9 58 06.32 +15 58 19.7  _{”}
4.23674 9 57 55.23 +15 55 53.1 17.8 Offutt
4.26837 9 57 54.05 +15 55 38.8 17.8  _{”}
4.48063 9 57 46.28 +15 54 00.2 17.7 Balam
4.48720 9 57 46.07 +15 53 55.3 17.5  _{”}
4.60091 9 57 42.51 +15 53 05.5 17.3 Kojima
4.61927 9 57 41.84 +15 52 58.1 17.1  _{”}

T. Kobayashi (Oizumi). 0.41-m f/4.3 reflector. H. Abe (Yatsuka). 0.26-m f/6 reflector. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector. Measurer Tichý.

Unfiltered images show suggestion of 10′′ tail in p.a. 300^{o}. W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien. Head slightly diffuse;

definite, very faint tail in p.a. 284^{o}, 17′′ long on Feb. 3, 8′′ on Feb. 4. T. Kojima (Chiyoda). 0.25-m f/6 reflector. P. Pravec and L. Šarounová (Ondřejov). 0.65-m f/3.6 reflector. D. D. Balam (Victoria). 0.5-m reflector.

1997 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden